OpenStudy (anonymous):

I need some help on these problems: 1) A swimming pool measures 9.0 m long by 3.5 m wide by 1.75 m deep. What mass of water in metric tons (1 metric ton = 1000 kg) does the pool contain when filled? The density of the water is 0.991 g/cm3. 2) If a liquid has a density of 1.17 g/cm3, how many liters of the liquid have a mass of 3.75 kg? 3) When a sample of a metal alloy that has the mass of 9.65 g is placed into a graduated cylinder containing water, the volume reading in the cylinder increases from 16.0mL to 19.5mL. What is the density of the alloy sample in grams per cubic meter?

OpenStudy (anonymous):

number 2 should be M=3.75/1.17=3.2 Liters, so the volume of the liquid is 3.2 Liters.

OpenStudy (anonymous):

3. the amount of water displaced is = to the volume of the object so you have 3.5 ml displaced. therefore, the volume of the object is 3.5 ml or 3.5 cm^3 density is mass / volume 9.65g / 3.5cm^3 = 2.8 g/cm^3

OpenStudy (anonymous):

Yayy! That means I number 2 right! Thanks! :)

OpenStudy (anonymous):

no problem:)

OpenStudy (anonymous):

I looked at number one and I'm like what the hell is this? :/

OpenStudy (anonymous):

i don't even know! teachers are killing us i swear